OpenStudy (anonymous):

Write two equations in standard form that are equivalent to the given equation. 1.5x+10y=15 2.-9x-12y=6 3.-2x+4y=-5

OpenStudy (saifoo.khan):

Maxxy!

OpenStudy (anonymous):

Saifooooo

OpenStudy (saifoo.khan):

do u want them y = mx + b form?

OpenStudy (anonymous):

No in standard form Ax+By=C

OpenStudy (anonymous):

Just write two other equations that are same and in standard form

OpenStudy (saifoo.khan):

5x+10y=15 x + 2y = 3, 10x+20y = 30

OpenStudy (anonymous):

huh

OpenStudy (anonymous):

Ohhhh

OpenStudy (anonymous):

ok can you do other 2

OpenStudy (saifoo.khan):

-9x-12y=6 3x -4y = 2, -18x -24y = 12

OpenStudy (saifoo.khan):

Now try yourself the third one! ;)

OpenStudy (anonymous):

k

OpenStudy (saifoo.khan):

when you solve it, write them here i will give them a chk.

OpenStudy (anonymous):

k

OpenStudy (anonymous):

k i have no idea how